Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson

Clear, seeing .4-.5"

7:30 pm (2:30 UT) start 2-hours of H-band science observations of MACS 0717 (Rex)

Still beautiful seeing .5"

9:30 pm (4:30 UT) start telluric acquisition. Error when moving mask into FPU.

"Error while trying to move mask to the focal plane unit at MOS unit 'rmi://localhost:1098/ONE_MOSUnit'!"

"Error setting the MOS position."

"error while executing transition:'focal plane unit clamp x - move motor until state is reached'"

Dave explains this means the mask is currently only being held in place by one of the clamps. Dave is calling Germany.

1:00 am (8:00 UT) Peter has just gotten back to us and fixed the problem in a few minutes. We lost 3.5 hours waiting to hear from Germany.

Returning to do telluric standard for Rex program-- more MOS errors.

Peter fixed the problem.

1:30 am (8:30 UT) Starting GOODS (GN26,Rujopakarn) observation.

seeing is 0.25"!

2:10 am (9:10 UT) seeing is 0.4.

3:00 am (10:00 UT) Starting AEGIS 1418 (Rujopakarn) observation.

seeing is 0.3"

4:30 am (11:30 UT) Starting Abell 2218b (Rujopakarn) observation.

6:00am (13:00 UT) stopped wavefront sensing, doing calibrations and darks.
Topic revision: r6 - 11 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback